# Multimodal ArXiv: A Dataset for Improving Scientific Comprehension of Large Vision-Language Models

### v2

@@ Abstract
Large vision-language models (LVLMs), exemplified by GPT-4V, excel across diverse tasks involving concrete images from natural scenes. However, their ability to interpret abstract figures, such as geometry shapes and scientific plots, remains limited due to a scarcity of training datasets in scientific domains. To fill this gap, we introduce Multimodal ArXiv, consisting of ArXivCap and ArXivQA, for enhancing LVLMs scientific comprehension. ArXivCap is a figure-caption dataset comprising 6.4M images and 3.9M captions sourced from 572K ArXiv papers spanning various scientific domains. Drawing from ArXivCap, we introduce ArXivQA, a question-answering dataset generated by prompting GPT-4V based on scientific figures. ArXivQA greatly enhances LVLMs’ mathematical reasoning capabilities, achieving a 10.4% absolute accuracy gain on a multimodal mathematical reasoning benchmark. Furthermore, employing ArXivCap, we devise four vision-to-text tasks for benchmarking LVLMs. Evaluation results with state-of-the-art LVLMs underscore their struggle with the nuanced semantics of academic figures, with domain-specific training yielding substantial performance gains. Our error analysis uncovers misinterpretations of visual context, recognition errors, and the production of overly simplified captions by current LVLMs, shedding light on future improvements. [^1] Project page: https://mm-arxiv.github.io.
@@
Lei Lithanks: Equal ${}^{\dagger}$, Yuqi $*{}^{*}$${}^{\dagger}$, Runxin ${}^{\ddagger}$, Peiyi ${}^{\ddagger}$ Xiachong ${}^{\dagger}$, Lingpeng ${}^{\dagger}$, Qi ${}^{\dagger}$$^{\dagger}$he University of Hong Kong ${}^{\ddagger}$eking University {nlp.lilei, runxinxu, wangpeiyi9979, xiachongfeng1996}@gmail.com wangyuqi@connect.hku.hk{lpk, liuqi}@cs.hku.hk
@@ 1 Introduction
Large vision-language models (LVLMs), which integrate large language models (LLMs) (Brown et al., 2020a; Touvron et al., 2023) with pre-trained vision encoders through cross-modal alignment training (Madureira, 2021; Liu et al., 2023b; Li et al., 2023d), have demonstrated remarkable perceptual and cognitive capabilities in processing concrete images from everyday scenes (OpenAI, 2023; Fu et al., 2023; Yang et al., 2023a). However, recent studies have exposed their limited comprehension of abstract figures, such as geometric shapes in multimodal mathematical reasoning (Lu et al., 2023) and scientific plots (Yue et al., 2023). The inadequacy of training datasets in scientific domains that involve complex reasoning with abstract figures is the main underlying cause.
@@
We validate the effectiveness of our Multimodal ArXiv from two dimensions: reasoning ability measured by QA accuracy and generation performance through captioning-like tasks. Our experiments demonstrate that ArXivQA brings a significant 10.4% absolute accuracy boost on the MathVista (Lu et al., 2023), a challenging benchmark for multimodal mathematical reasoning. Additionally, detailed analysis uncovers the relationship between paper domains and fine-grained task performance. Moreover, using ArXivCap, we define four generation tasks of varying complexity to benchmark the ability of LVLMs to comprehend scientific plots: (1) captioning a single academic figure, (2) generating overall summaries for multiple sub-figures, (3) in-context image captioning given previous figure-caption pairs, and (4) generating paper titles from figure-caption pairs. We examine various LVLMs, including open-source models as well as proprietary models including GPT-4V (OpenAI, 2023) and Bard (Google, 2023). Evaluation results reveal that despite that current LVLMs still face challenges generating faithful captions for scientific figures, in-domain training on our dataset yields substantial performance improvements across all four tasks. Manual error analysis underscores that LVLMs still suffer from misinterpretation of the visual context, recognition errors, and overly simplified captions, paving the way for future studies.
@@ 2 Related Work
Recent advancements in LVLMs have seen notable progress in model architecture, training paradigms, and dataset creation.
Model Architecture LVLMs typically comprise three core modules: (i) a vision encoder for image feature extraction, (ii) a modality alignment module to integrate visual features into the language model embedding space, and (iii) an LLM backbone for decoding multimodal context. CLIP (Radford et al., 2021) is widely used for image encoding, while LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023) serve as popular choices for LLMs. The alignment module varies from simple linear projections (Liu et al., 2023b; Zhu et al., 2023) to more complex architectures like gated cross-attention layers  substantiated by Flamingo and IDEFICS(Alayrac et al., 2022; Awadalla et al., 2023). Innovations such as the Q-Former module in BLIP2 (Li et al., 2023b) and instruction integration in InstructBLIP (Dai et al., 2023) further enhance alignment capabilities. Additionally, Fuyu-8B (Bavishi et al., 2023) introduces a novel framework mapping raw image pixels directly to the LLM embedding space.
Training Paradigms PaLI-X (Chen et al., 2023b) investigates scaling effects of both vision encoders and language models, highlighting the advantages of scaling both components. Qwen-VL (Bai et al., 2023b) increases input image resolution and explores different module unfreezing strategies. Alignment methodologies such as RLHF training (Ouyang et al., 2022), e.g., LLaVA-RLHF (Sun et al., 2023), and preference optimization through AI feedback (Li et al., 2023c) demonstrate improved alignment with human preferences.
[…]
@@ 3.1 ArXivCap
Paper Filtering with Publication Type ArXivCap is extracted from ArXiv Clement et al. (2019), which is under CC-0 licence for modification and distribution. The raw files of papers posted on ArXiv tar files before June 2023 are downloaded. To ensure the quality of our dataset, we employ a rigorous selection process to filter potentially low-quality papers that might influence the figure-caption pair quality. Firstly, we retrieve meta-information for papers from Semantic Scholar Kinney et al. (2023), which contains the publication record for each paper. Papers with publication types JournalArticle, Conference, or Review are kept as we assume the peer-review process could ensure the overall figure-caption quality is satisfactory. We further treat papers with titles longer than 100 words, or with an abstract longer than 300 words according to common submission requirements and remove them.
Figure-Caption Pair Extraction Images and captions are extracted from the original LaTeX files by matching the syntax. We further use a robust tool ImageMagisk ImageMagick Studio LLC to convert images into JPEG format for easy processing. The extracted images and captions are stored in a designed chunk structure, which consists of either a single figure-caption pair or multiple figures with their respective sub-captions and a main caption for the overall description. This format is more consistent with the layout of academic papers, and Figure 2 illustrates the chunk structure.
@@
Caption Cleaning and Image Filtering After a manual inspection of the initially collected dataset, we design several transformations to clean the captions and filter the images. Caption Cleaning: (i) Chunks with captions shorter than 5 words are removed; (ii) For captions with LaTex expressions such as math formulas and references, we apply the pylatexenc to transform the LaTeX to text with math formulas retained, citations to a special symbol <cit.>, references to <ref>. An illustration of caption cleaning can be found in Appendix A.1. Image Filtering: We remove images that are deemed to be problematic according to the following rules: (i) Images with an aspect ratio larger than 100; (ii) Images with the shortest edge shorter than 224 pixels; and (iii) Images with pixel numbers larger than the decompression bombs threshold.
After these processes, 100 pairs are sampled to perform an additional manual inspection, where we found all of these pairs contained clear images and correct caption descriptions. We provide visualized samples in Appendix A.3.
3.1.1 Statistics of ArXivCap
@@
Table 1 lists the dataset statistics. ArXivCap consists of 572K papers, containing 6.4M high-quality images in total with 193M words. A word cloud illustration of captions can be found in the Appendix A.2. Figure 3 demonstrates the paper domain distribution extracted from ArXiv, where we find that our ArXivCap covers 32 paper domains, such as computer science, mathematics, physics, and economics. As shown in Table 2, compared with previous scientific figure datasets, our ArXivCap is the largest figure-caption dataset collected from real papers and covers a wide range of scientific domains, serving as a valuable resource for improving and benchmarking LVLMs.
@@ 3.2 ArXivQA
As our ArXivCap contains diverse images from scientific domains, we assume that learning to answer questions about these figures could boost scientific reasoning ability. Following the successful practice of LLaVA (Liu et al., 2023b), we adopt GPT-4V to generate instruction-tuning datasets for generating the QA pairs based on the figures extracted from scientific papers. Specifically, we design a prompting template to query GPT-4V for generating QA pairs based on 35K images randomly sampled from our ArXivCap. Table 9 in Appendix A.4 provides the template we used for the prompt. The generated pairs are parsed according to the format requirement and we discard the samples without options and rationales. There are 100K QA pairs after filtering the invalid samples. The dataset comprises questions with an average word count of 16.98 for the question text, and options for the question with an average word count of 31.86. On average, there are 4.20 options per question. Appendix A.3 provides samples from the ArXivQA dataset.
@@ 4 Experiments
In this section, we conduct experiments to (i) validate the effectiveness of ArXivQA (§4.1) and (ii) benchmark the LVLMs capability to comprehend scientific figures with our ArXivCap (§4.2).
@@ 4.1.1 Experimental Settings
We adopt Qwen-VL-Chat (Bai et al., 2023b) as the backbone, and fine-tune it on our ArXivCap and ArXivQA datasets for three epochs with a learning rate of 1e-5 following the original paper. Models are evaluated on MathVista (Lu et al., 2023), a benchmark that requires fine-grained, deep visual understanding and compositional reasoning. MathVista contains 6,141 examples, consisting of five multimodal tasks Figure QA, Geometry Problem Solving, Math word problem, Text Book QA, and Visual QA. We select the multiple-choice questions in the benchmark to avoid the inconsistency of answer parsing. We compute the accuracy scores on the testmini set and adopt the provided prediction files for calculating the baseline performance.
@@ 4.1.2 Results
${}^{\dagger}$ 41.4 22.0 18.2 34.6 44.6 33.7
${}^{\dagger}$ 41.4 19.9 45.5 45.8 57.6 39.3
${}^{\dagger}$ 44.0 26.7 40.9 45.8 44.6 39.3
Qwen-VL-Chat 48.3 19.1 22.7 46.7 57.6 40.0
 w/ ArXivCap 39.7 19.8 27.2 39.7 52.1 36.2
 w/ ArXivQA 44.8 34.0 27.3 70.0 64.1 50.2
 w/ ArXivQA + ArXivCap 44.0 37.6 27.3 68.2 63.0 50.4
${}^{\dagger}$ 38.8 51.1 27.3 64.5 51.1 50.0
${}^{\dagger}$ 52.6 51.8 54.5 83.2 66.3 61.9
Table 3: Evaluation on MathVista dataset. ArXivCap and ArXivQA together enhance Qwen-VL-Chat’s overall performance, surpassing that of the commercial model Bard. ${}^{\dagger}$ denotes results based on the original predictions from Lu et al. (2023). The best results are highlighted in bold, while the second-best scores are marked with underline.
As shown in Table 3, fine-tuning on our Multimodal ArXiv, especially on the ArXivQA dataset, consistently boosts the performance, helping the open-sourced Qwen-VL-Chat achieve a comparable overall MathVista reasoning performance. Due to the wide coverage of the scientific figures, the performance gain mainly comes from significantly improved Geometry Problem Solving, Textbook QA, and Visual QA tasks. For example, after fine-tuning on the ArXivQA dataset, the accuracy is increased from 46.7% to 70.0% and from 19.1% to 34.0% on Geometry Problem Solving and Textbook QA tasks, respectively. The improvement on Math Word Problem is marginal, where we think the domain-specific data augmentation can be further explored with a curated filtering dataset on our dataset (Gao et al., 2023). On the contrary, the accuracy of Figure QA deteriorates slightly compared with the original backbone model, which we attribute to the fact that most of the plots in the Figure QA evaluation are sampled from synthesized datasets such as DVQA (Kafle et al., 2018), exhibiting great gaps between real-world paper figures.

### v3

@@ Abstract
Large vision-language models (LVLMs) excel across diverse tasks involving concrete images from natural scenes. However, their ability to interpret abstract figures, such as geometry shapes and scientific plots, remains limited due to a scarcity of training datasets in scientific domains. To fill this gap, we introduce Multimodal ArXiv, consisting of ArXivCap and ArXivQA, for enhancing LVLMs scientific comprehension. ArXivCap is a figure-caption dataset comprising 6.4M images and 3.9M captions, sourced from 572K ArXiv papers spanning various scientific domains. Drawing from ArXivCap, we introduce ArXivQA, a question-answering dataset generated by prompting GPT-4V based on scientific figures. ArXivQA greatly enhances open-sourced LVLMs’ mathematical reasoning capabilities, achieving a 10.4% absolute accuracy gain on a multimodal mathematical reasoning benchmark. Furthermore, employing ArXivCap, we devise four vision-to-text tasks for benchmarking LVLMs. Evaluation results with state-of-the-art LVLMs underscore their struggle with the nuanced semantics of academic figures, while domain-specific training yields substantial performance gains. Our error analysis uncovers misinterpretations of visual context, recognition errors, and the production of overly simplified captions by current LVLMs, shedding light on future improvements. [^1] Datasets and models are released at our project page: https://mm-arxiv.github.io.
@@
Lei Lithanks: Equal Contribution., Yuqi Wang, Runxin Xu, Peiyi Wang Xiachong Feng, Lingpeng Kong, Qi Liu The University of Hong Kong Peking University {nlp.lilei, runxinxu, wangpeiyi9979, xiachongfeng1996}@gmail.com wangyuqi@connect.hku.hk{lpk, liuqi}@cs.hku.hk
@@ 1 Introduction
Large vision-language models (LVLMs), which integrate large language models (LLMs) (Brown et al., 2020a; Touvron et al., 2023) with pre-trained vision encoders through cross-modal alignment training (Madureira, 2021; Liu et al., 2023b; Li et al., 2023d), have demonstrated remarkable perceptual and cognitive capabilities in processing concrete images from everyday scenes (OpenAI, 2023; Fu et al., 2023; Yang et al., 2023a; Reka, 2024). However, recent studies have shown that open-source LVLMs struggle to understand abstract figures, such as geometric shapes in multimodal mathematical reasoning (Lu et al., 2023; Zhang et al., 2024b) and scientific plots (Yue et al., 2023). The inadequacy of training datasets in scientific domains that involve complex reasoning with abstract figures is the main underlying cause.
@@
We validate the effectiveness of our Multimodal ArXiv dataset from two dimensions: reasoning ability measured by QA accuracy and generation performance through novel vision-to-text tasks. Our experiments demonstrate that ArXivQA brings a significant 10.4% absolute accuracy boost for Qwen-VL-Chat (Bai et al., 2023b), on the MathVista (Lu et al., 2023), a challenging benchmark for multimodal mathematical reasoning. Additionally, detailed analysis uncovers the relationship between paper domains and fine-grained task performance. Moreover, using ArXivCap, we define four generation tasks of varying complexity to benchmark the ability of LVLMs to comprehend scientific plots: (1) captioning a single academic figure, (2) generating overall summaries for multiple sub-figures, (3) in-context figure captioning given previous figure-caption pairs, and (4) generating paper titles from figure-caption pairs. We examine various LVLMs, including open-source models as well as proprietary models including GPT-4V (OpenAI, 2023) and Gemini 1.0 Pro Vision (Gemini Team, 2023). Evaluation results reveal that despite that current LVLMs still face challenges generating faithful captions for scientific figures, in-domain training on our dataset yields substantial performance improvements across all four tasks. Manual error analysis underscores that LVLMs still suffer from misinterpretation of the visual context, recognition errors, and overly simplified captions, paving the way for future studies.
@@ 2 Related Work
Recent advancements in LVLMs have seen notable progress in model architecture, training paradigms, and dataset creation (Zhang et al., 2024a).
Model Architecture
LVLMs typically comprise three core modules: (i) a vision encoder for image feature extraction, (ii) a modality alignment module to integrate visual features into the language model embedding space, and (iii) an LLM backbone for decoding multimodal context. CLIP (Radford et al., 2021) is widely used for image encoding, while LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023) serve as popular choices for LLMs. The alignment module varies from simple linear projections (Liu et al., 2023b; Zhu et al., 2023) to more complex architectures like gated cross-attention layers  substantiated by Flamingo and IDEFICS (Alayrac et al., 2022; Awadalla et al., 2023). Innovations such as the Q-Former module in BLIP2 (Li et al., 2023b) and instruction integration in InstructBLIP (Dai et al., 2023) further enhance alignment capabilities. Additionally, Fuyu-8B (Bavishi et al., 2023) introduces a novel framework mapping raw image pixels directly to the LLM embedding space.
Training Paradigms
Regarding the training recipes, PaLI-X (Chen et al., 2023b) investigates the scaling effects of both vision encoders and language models, highlighting the advantages of scaling both components. Qwen-VL (Bai et al., 2023b) increases input image resolution and explores different module unfreezing strategies. Alignment methodologies such as RLHF training (Ouyang et al., 2022), e.g., LLaVA-RLHF (Sun et al., 2023), and preference optimization through AI feedback (Li et al., 2023c) demonstrate effectiveness in aligning LVLMs with human preferences.
Dataset Curation
Dataset quality significantly impacts LVLM performance. Modality alignment training often utilizes web-scale image-caption pairs such as Laion-400M (Schuhmann et al., 2021), with recent studies favoring cleaned captions (Chen et al., 2023a; Yu et al., 2023). Instruction fine-tuning (IFT) helps LVLMs respond according to user queries, triggering the exploration of high-quality IFT datasets. Efforts include multimodal instruction collections such as MultiInstruct (Xu et al., 2023) and M3IT (Li et al., 2023d), dialog-style datasets such as LLaVA (Liu et al., 2023b) and domain-specific datasets for medical (Li et al., 2023a) and text-rich images (Zhang et al., 2023). In the scientific domain, FigCAP (Chen et al., 2019) and FigureQA (Kahou et al., 2017) are created based on synthetic figures. DVQA (Kafle et al., 2018) creates heuristic-based questions for bar charts only. SciCap (Hsu et al., 2021), SciCap+ (Yang et al., 2023b), and M-Paper (Hu et al., 2023) collect figure-caption pairs from specific domains such as computer science. Compared with these datasets, our ArXivCap is sourced from diverse scientific domains with a much larger scale, enabling more comprehensive improvements and evaluations. Besides, we employ GPT-4V for creating ArXivQA with challenging questions, showcasing its effectiveness in boosting the mathematical reasoning ability of LVLMs.
@@ 3.1 ArXivCap
Paper Filtering with Publication Type: ArXivCap is extracted from ArXiv Clement et al. (2019), which is under CC-0 licence for modification and distribution. The raw files of papers posted on ArXiv tar files before June 2023 are downloaded. To ensure the quality of our dataset, we employ a rigorous selection process to filter potentially low-quality papers that might influence the figure-caption pair quality. Firstly, we retrieve meta-information for papers from Semantic Scholar Kinney et al. (2023), which contains the publication record for each paper. Papers with publication types JournalArticle, Conference, or Review are kept as we assume the peer-review process could ensure the overall figure-caption quality is satisfactory. We further exclude papers with titles exceeding 100 words or abstracts longer than 300 words, in alignment with common submission requirements.
Figure-Caption Pair Extraction: Images and captions are extracted from the original LaTeX files by matching the syntax. We further use a robust tool ImageMagisk ImageMagick Studio LLC to convert images into JPEG format for easy processing. The extracted images and captions are stored in a designed chunk structure, which consists of either a single figure-caption pair or multiple figures with their respective sub-captions and a main caption for the overall description. This format is more consistent with the layout of academic papers, and Figure 2 illustrates the chunk structure.
@@
Caption Cleaning and Image Filtering: After a manual inspection of the initially collected dataset, we design several transformations to clean the captions and filter the images.
Caption Cleaning: (i) Chunks with captions shorter than 5 words are removed; (ii) For captions with LaTeX expressions such as math formulas and references, we apply the pylatexenc [^2] https://github.com/phfaist/pylatexenc to transform the LaTeX to text with math formulas retained, citations to a special symbol <cit.>, references to <ref>. An illustration of caption cleaning can be found in Appendix A.1.
Image Filtering: We remove images that are deemed to be problematic according to the following rules: (i) Images with an aspect ratio larger than 100; (ii) Images with the shortest edge shorter than 224 pixels; and (iii) Images with pixel numbers larger than the decompression bombs threshold.
After these processes, 100 pairs are sampled to perform an additional manual inspection, where we found all of these pairs contained clear images and correct caption descriptions. We provide visualized figure-caption pairs in Appendix A.2.
Statistics of ArXivCap
@@
Table 1 lists the dataset statistics. ArXivCap consists of 572K papers, containing 6.4M high-quality images in total with 193M words. A word cloud illustration of captions can be found in the Appendix A.3. Figure 3 demonstrates the paper domain distribution extracted from ArXiv, where we find that our ArXivCap covers 32 domains, such as computer science, mathematics, physics, and economics. As shown in Table 2, compared with previous scientific figure datasets, our ArXivCap is the largest figure-caption dataset collected from real papers and covers a wide range of scientific domains, serving as a valuable resource for improving and benchmarking LVLMs.
@@ 3.2 ArXivQA
As our ArXivCap contains diverse images from scientific domains, we assume that learning to answer questions about these figures could boost scientific reasoning ability. Following the successful practice of LLaVA (Liu et al., 2023b), we adopt GPT-4V to generate instruction-tuning datasets for generating the QA pairs based on the figures extracted from scientific papers. Specifically, we design a prompting template to query GPT-4V for generating QA pairs based on 35K images randomly sampled from our ArXivCap. Table 11 in Appendix A.4 provides the template we used for the prompt. The generated pairs are parsed according to the format requirement and we discard the samples without options and rationales. There are 100K QA pairs after filtering the invalid samples. The dataset comprises questions with an average word count of 16.98 for the question text. On average, there are 4.20 options per question and the average length of the text for a single option is 7.59 words. Appendix A.2 provides samples from the ArXivQA dataset.
As a preliminary study, we sample 1,000 samples from ArXivQA and prompt open-sourced LVLMs to predict answers given the questions and options. A simple prompt is designed to employ GPT-4 for extracting the answer label from the model generations. For human performance, we ask four authors to perform predictions on a 100-sample subset (where 17 samples are from the CS domain). Each of them is asked to answer 50 samples and the accuracy scores are obtained by averaging two annotators. As shown in Table 3, most models struggle to perform satisfactorily on the ArXivQA dataset, falling far behind human performance. This verifies our premise that current open-sourced LVLMs fail to understand scientific figures. We also notice that simply increasing the model scale from 7B (LLaVa-1.5-7B) to 13B (LLaVa-1.5-13B) does not yield a significant boost, which indicates that the ability for multimodal mathematical reasoning cannot be simply acquired from the LLM-side only.
Model Accuracy
InstructBLIP-Vicuna7B 7.0%
LLaVA-1.5-7B 44.2%
LLaVA-1.5-13B 46.8%
OpenFlamingo-9B 9.9%
IDEFICS-Instruct-9B 34.5%
Qwen-VL-Chat 46.6%
Human (100-sample Subset) 80.0%
Human (CS subset) 88.2%
Table 3: Evaluation results on the sampled 1,000 ArXivQA samples.
IDEFICS-Instruct-9B 41.4 22.0 18.2 34.6 44.6 33.7
InstructBLIP-Vicuna13B 41.4 19.9 45.5 45.8 57.6 39.3
LLaVa-v1.5-13B 44.0 26.7 40.9 45.8 44.6 39.3
Qwen-VL-Chat-7B 48.3 19.1 22.7 46.7 57.6 40.0
[…]
Qwen-VL-Chat-7BArXivCap + ${}_{\text{ArXivCap + ArXivQA}}$ 44.0 37.6 27.3 68.2 63.0 50.4
Bard 38.8 51.1 27.3 64.5 51.1 50.0
GPT-4V 52.6 51.8 54.5 83.2 66.3 61.9
Table 4: Evaluation on MathVista dataset. ArXivCap and ArXivQA together enhance Qwen-VL-Chat’s overall performance, surpassing that of the commercial model Bard. denotes results based on the original predictions from Lu et al. (2023). The best results are highlighted in bold, while the second-best scores are marked with underline.
@@ 4 Experiments
We conduct experiments to (i) validate the effectiveness of ArXivQA for boosting multimodal scientific reasoning for open-source LVLMs (§4.1) and (ii) benchmark LVLMs capability to comprehend scientific figures with ArXivCap (§4.2).
@@ 4.1.1 Experimental Settings
We adopt Qwen-VL-Chat-7B (Bai et al., 2023b) as the backbone due to its support for interleaved image-text input formats and high-resolution images. We fine-tune it on our ${}_{\text{ArXivCap}}$), ${}_{\text{ArXivQA}}$) and combination of these two datasets (Qwen-VL-Chat-7BArXivCap + ${}_{\text{ArXivCap + ArXivQA}}$) for three epochs with a learning rate of 1e-5 following the original paper. We combine the answer and the rationale in ArXivQA to form the target output during training. Models are evaluated on MathVista (Lu et al., 2023), a benchmark that requires fine-grained, deep visual understanding and compositional reasoning. MathVista contains 6,141 examples, consisting of five multimodal tasks Figure QA, Geometry Problem Solving, Math word problem, Text Book QA, and Visual QA. We select 478 multiple-choice questions in the testmini split to avoid the inconsistency of answer parsing. We compute the accuracy scores and adopt the provided prediction files for calculating the baseline performance.
@@ 4.1.2 Results
As shown in Table 4, fine-tuning on our Multimodal ArXiv, especially on the ArXivQA dataset, consistently boosts the performance, helping the open-sourced Qwen-VL-Chat achieve a comparable overall MathVista reasoning performance. Due to the wide coverage of the scientific figures, the performance gain mainly comes from significantly improved Geometry Problem Solving, Textbook QA, and Visual QA tasks. For example, after fine-tuning on the ArXivQA dataset, the accuracy is increased from 19.1% to 34.0% and from 46.7% to 70.0% on Geometry Problem Solving and Textbook QA tasks, respectively. The improvement on Math Word Problem is marginal, where we think the domain-specific data augmentation can be further explored with a curated filtering dataset on our dataset (Gao et al., 2023). On the contrary, the accuracy of Figure QA deteriorates slightly compared with the original backbone model, which we attribute to the fact that most of the plots in the Figure QA evaluation are sampled from synthesized datasets such as DVQA (Kafle et al., 2018), exhibiting great gaps between real-world paper figures.